Plot graph of

Qd=20-4p, where p=#2,4

Qs=12+3p, where p=#2,4


1
Expert's answer
2022-01-20T13:48:44-0500

In equilibrium Qd = Qs, so:

20 - 4p = 12 + 3p,

p = 8/7.

So, both prices p = 2 and p = 4 are above equilibrium, where Qd = Qs.